OpenStudy (anonymous):

Find the equation of the plane through (1,1,1) and containing the line which is the intersection of the planes x-y=2 and y-z=1

OpenStudy (anonymous):

Can someone explaine to me how this is done

OpenStudy (amistre64):

the line of intersecting planes is the cross of their direction vectors

OpenStudy (anonymous):

First of all you need to understand how a plane is formed

OpenStudy (amistre64):

<1,-1,0> x <0,1,-1> = intersection vector

OpenStudy (anonymous):

How did you get that @amistre64?

OpenStudy (amistre64):

i pulled of the normals from the plane eqs and crossed them

OpenStudy (amistre64):



OpenStudy (anonymous):

okay, so then you did the cross product

OpenStudy (amistre64):

yes, cross the normals of the planes given to find the line asked for; then we work from that thing

OpenStudy (anonymous):

okay, so the cross product gives us the vector: i+j+k

OpenStudy (amistre64):

should i trust you for that? :)

OpenStudy (amistre64):

the rest of the question is vague or mistyped

OpenStudy (amistre64):

we need to find 2 points on this intersecting line tho

OpenStudy (anonymous):

nope that is how it is written on my sample test

OpenStudy (amistre64):

find a point that both intersecting planes have in common

OpenStudy (amistre64):

x-y=2 and y-z=1 when x=0; y=-2 when y=-2; z=-3?

OpenStudy (amistre64):

so the point (0,-2,-3) is on our intersecting line we can use that and another point on the line to create 2 vectors to our (1,1,1) point

OpenStudy (amistre64):

line: x= 0+t y=-2+t z=-3+t pick your favorite t

OpenStudy (anonymous):

h

OpenStudy (amistre64):

h aint a number that we can define a point with

OpenStudy (anonymous):

oh i meant 6

OpenStudy (amistre64):